OpenStudy (anonymous):

Suppose you choose a marble from a bag containing 2 red marbles, 5 white marbles, and 3 blue marbles. You return the first marble to the bag and then choose again. Find P(red and blue).

OpenStudy (across):

2 : red 5 : white + 3 : blue ------- 10 : total P(red)=1/5 P(white)=1/2 P(blue)=3/10 P(red and blue)=P(red)P(blue)=(1/5)(3/10)=3/50

OpenStudy (anonymous):

are you looking for the P. for one red and one blue?